OpenStudy (anonymous):

Given 2 points, find the slope using the slope formula. (12,5) and (9,8)

OpenStudy (karatechopper):

Do you know the slope formula?

OpenStudy (anonymous):

Yeah, I know it.

OpenStudy (karatechopper):

Ok good! Then can you tell me what it is?

OpenStudy (anonymous):

\[slope = (DeltaY)/(DeltaX)\] ;then the slope =(Y2-Y1)/(X2-X1) then the slope =(8-5)/(9-12)=(3)/(-3) ; so the slope =-1

OpenStudy (anonymous):

Thank you. :)